# WenetSpeech4TTS: A 12,800-hour Mandarin TTS Corpus for Large Speech Generation Model Benchmark

### v1

@@ 2.6 Quality filtering
After the preceding steps, we conduct a quality-based data filtration. Segments with multiple speakers, as well as those with speech recognition confidence scores below $0.7$, are excluded. The refined dataset, designated as the WenetSpeech4TTS corpus, comprises $12,800$ hours of speech segments. Some interval between segments, which are not included in WenetSpeech, are merged into WenetSpeech4TTS through the merging operation. This enables us to obtain a larger dataset.
@@ 4 Experiments
To validate the usability of the WenetSpeech4TTS dataset and establish baselines on benchmark for large-scale TTS models, VALL-E [2] and NaturalSpeech 2 [4] are trained on the Basic set and finetuned on the other two sets with higher quality successively.
@@ 4.1 VALL-E and NaturalSpeech 2
We have replicated and modified VALL-E, one of the most popular large-scale TTS models. In addition, we have utilized the Amphion [^6] https://github.com/open-mmlab/Amphion toolkit [18] toolkit to implement NaturalSpeech 2.
@@
The model configurations of NaturalSpeech 2 follow the default configurations in the Amphion toolkit. The data preprocessing process is consistent with VALL-E, except that we use the default Encodec as the acoustic codec of NaturalSpeech 2. Additionally, we sample $2,000$ hours of labeled data from WenetSpeech4TTS Basic to train the ASR model in Kaldi  [^9] https://github.com/kaldi-asr/kaldi toolkit, and then obtain external duration information of all labeled data for NaturalSpeech 2 training. We train the model for a total of $600$k steps on $8$ NVIDIA A6000 GPUs. The training and fine-tuning strategy of NaturalSpeech 2 is similar to that of VALL-E.

### v2

@@ 2.6 Quality filtering
After the preceding steps, we conduct a quality-based data filtration. Segments with multiple speakers, as well as those with speech recognition confidence scores below $0.7$, are excluded. The refined dataset, designated as the WenetSpeech4TTS corpus, comprises $12,800$ hours of speech segments. Some intervals between segments, which are not included in WenetSpeech, are merged into WenetSpeech4TTS through the merging operation. This enables us to obtain a larger dataset.
@@ 4 Experiments
To validate the usability of the WenetSpeech4TTS dataset and establish baselines on benchmark for large-scale TTS models, two popular large-scale TTS models, VALL-E [2] and NaturalSpeech 2 [4], are trained on the Basic set and finetuned on the other two sets with higher quality successively. We open-source these model weights along with our dataset.
@@ 4.1 VALL-E and NaturalSpeech 2
The model configurations of NaturalSpeech 2 follow the default configurations in the Amphion toolkit [^8] https://github.com/open-mmlab/Amphion  [20]. The data preprocessing process is consistent with VALL-E, except that we use the default Encodec as the acoustic codec of NaturalSpeech 2. Additionally, we sample $2,000$ hours of labeled data from WenetSpeech4TTS Basic to train the ASR model in Kaldi  [^9] https://github.com/kaldi-asr/kaldi toolkit, and then obtain external duration information of all labeled data for NaturalSpeech 2 training. We train the model for a total of $600$k steps on $8$ NVIDIA A6000 GPUs. The training and fine-tuning strategy of NaturalSpeech 2 is similar to that of VALL-E.